OpenStudy (anonymous):

Help please??? Solve the following system of equations using matrices. 2x + 4y = 16 -4x – 8y = 32 A. (2, 2) B. (-2, -2) C. Consistent D. Inconsistent

OpenStudy (solomonzelman):

i would just use elimination after dividing the second equation by 2.

OpenStudy (solomonzelman):

or, you can divide the second equation by -2 and then tell me what clue do you have (it's ok if you don't come up with any thoughts).

OpenStudy (solomonzelman):

Do you have to (like just must) to use matrix to solve it?

OpenStudy (anonymous):

I think its inconsistent because it can't be A. or B.

OpenStudy (solomonzelman):

Can you tell me how many solutions it has?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 2x + 4y = 16 \\ -4x – 8y = 32 }\) divide the second equation by -2 and re-write the system of equations for me please.

OpenStudy (anonymous):

If I graph the equations they never cross so it would be inconsistent, right?

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

but you should obtain that result not by graphing

OpenStudy (solomonzelman):

you should have solved it. i would have helped you.

OpenStudy (solomonzelman):

i mean if you are still willing to go through it....

OpenStudy (anonymous):

I just wanted to check my answer to make sure I had done the problem correctly. Thank you for the help though <3